OpenStudy (anonymous):

If 7.65 grams of iron (III) oxide reacts with 6.85 grams of carbon monoxide to produce 5.10 g of pure iron, what are the theoretical yield and percent yield of this reaction? Be sure to show the work that you did to solve this problem. unbalanced equation: Fe2O3 + CO “yields”/ Fe + CO2

OpenStudy (aaronq):

So first balance the equation.

OpenStudy (anonymous):

okay can you give me a second to balance it because this one is a little harder than the last one

OpenStudy (aaronq):

take your time

OpenStudy (anonymous):

Okay I think it is Fe203 + 3CO "yields" 2Fe + 3CO2

OpenStudy (aaronq):

Its good! Now convert the masses given to moles. We need to determine which is the limiting reactant before continuing.

OpenStudy (anonymous):

okay just give a minute and again thanks for taking your time to help me

OpenStudy (aaronq):

yeah dude, no problem. you dont have to tell me to wait a minute, i know it takes some time.

OpenStudy (anonymous):

I know you start with the given Grams which is 7.65 but I forgot where to go from there

OpenStudy (anonymous):

I think then you need to add the elements masses together?

OpenStudy (aaronq):

divide the mass (m) by the molar mass (M) to get moles (n): \(n=\dfrac{m}{M}\)

OpenStudy (anonymous):

I am having trouble knowing what the molar mass is can you remind me?

OpenStudy (aaronq):

Add up the individual molar masses in the right proportions molar mass of iron = 55.8 g/mol molar mas of oxygen = 16 g/mol there are 2 Fe, 3 O, molar mass of \(Fe_2O_3\)= 2(55.8 g/mol)+3(16 g/mol)= 159.6 g/mol

OpenStudy (anonymous):

And just the mass is the one in the given right?

OpenStudy (aaronq):

yes 7.65 grams Fe2O3 you need to find the moles of CO as well

OpenStudy (anonymous):

CO is 58.9. and its that I am getting a big decimal number and I feel that it isn't right when I divide the mass by the molar mass

OpenStudy (anonymous):

I am getting 0.02976

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

CO is 12+16=28 g/mol

OpenStudy (aaronq):

so what are the moles of each?

OpenStudy (anonymous):

So why do I need to find the CO?

OpenStudy (aaronq):

Because you need to determine which of your starting products will run out first.

OpenStudy (anonymous):

So I am assuming I need to do the same to CO that I did to Fe2O3 so I got 0.27321

OpenStudy (aaronq):

Now to find the limiting reactant, we divide the moles of each reactant by their corresponding stoichiometric coefficient. And compare, the one with less moles is the limiting reactant.

OpenStudy (anonymous):

I think that the limiting reactant is either CO or Fe

OpenStudy (aaronq):

we need to know which one though.

OpenStudy (anonymous):

I think the limiting reactant is CO

OpenStudy (anonymous):

Am I right?

OpenStudy (aaronq):

sorry dude, the site is freezing on me

OpenStudy (aaronq):

what were your values?

OpenStudy (anonymous):

What do you mean "values"?

OpenStudy (aaronq):

".. divide the moles of each reactant by their corresponding stoichiometric coefficient."

OpenStudy (anonymous):

So does that mean I divide CO by the coefficient which is 3? and do the same for Fe?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

So I have 0.27321/3 which gives me 0.09107

OpenStudy (aaronq):

so whats the limiting reactant?

OpenStudy (anonymous):

and for Fe I think it is 0.04793/2 which gives me 0.023966. The reason I am typing everything I am doing is because I want to make sure I am doing it right.

OpenStudy (anonymous):

And the limiting reactant is the one with the lowest right?

OpenStudy (aaronq):

Okay you went ahead and finished the problem. Now find the mass of Fe with those moles

OpenStudy (aaronq):

sorry i thought we were a step behind

OpenStudy (anonymous):

Okay Now I am confused at which step we are at lol.

OpenStudy (aaronq):

You said "and for Fe.. " Did you find the moles of Fe? are you at that step? I thought we were still working with \(Fe_2O_3\)

OpenStudy (anonymous):

Oh yeah I did jumped ahead I am sorry I forgot we were doing reactants

OpenStudy (aaronq):

SO you have the moles of Fe? what is it?

OpenStudy (anonymous):

So Like you said Fe2O3 has 1 coefficient

OpenStudy (anonymous):

No I don't It's that I got confused in the middle of it sorry I am still at Fe2O3

OpenStudy (aaronq):

Okay, so we were finding the limiting reactant. Did you determine which had less moles after dividing by their coefficients?

OpenStudy (anonymous):

I think I did and I think the one with the less moles was Fe2O3 has less moles

OpenStudy (aaronq):

okay, good stuff. Now we make a ratio of the moles of the limiting reactant (\(Fe_2O_3\) in this case) and the moles of what we want to find (Fe in this case) with their coefficients (like in the previous problem).

OpenStudy (aaronq):

\(\sf \dfrac{moles~of ~Fe_2O_3}{Fe_2O_3's~ coefficient}=\dfrac{moles~of~Fe}{Fe's ~coefficient}\)

OpenStudy (anonymous):

Can you help me find the moles of Fe please?

OpenStudy (anonymous):

wait hold on

OpenStudy (aaronq):

i gotta get going. After you find the moles of Fe, you need to convert them to mass. (This is the theoretical yield.) Use: \(n=\dfrac{m}{M}\) After, to find the percent yield, use: \(\sf\% ~yield=\dfrac{actual~yield}{theoretical~yield}*100\%\)

OpenStudy (asib1214):